CHARACTERIZATION OF WEAKLY BOUND HCF_{3}–OCS, H_{2}CF_{2}–OCS, H_{3}CF–OCS, AND HCF_{3}–CO_{2} VAN DER WAALS COMPLEXES BY AB INITIO CALCULATIONS AND MICROWAVE SPECTROSCOPY.

\underline{MICHAL M. SERAFIN}, SEAN A. PEEBLES, Department of Chemistry, Eastern Illinois University, 600 Lincoln Ave., Charleston, IL 61920 USA.

High resolution Fourier-transform microwave spectroscopy has been utilized in a systematic study of the structural and dynamic properties of a series of fluorinated methane molecules complexed to carbonyl sulfide and carbon dioxide. Theoretical data from ab initio calculations at the MP2/6-311++G(2d, 2p) level have provided good agreement with the observed experimental values. The structures of HCF_{3}–OCS and H_{2}CF_{2}–OCS have been determined from experimentally obtained rotational spectra fit to a Watson A reduction Hamiltonian to within 4 kHz. The a-type rotational spectrum of H_{3}CF–OCS has been assigned, however, b-type transitions still need to be located. The HCF_{3}–CO_{2} spectrum showed a doubling of transitions into A and E states (by up to 1.5 MHz), and is currently fit to approximately 20 kHz using XIAM^{a}. The ab initio structures of all the complexes within the series are very similar with O⋯H distances consistently between 2.57-2.65 Å and O=C⋯C bond angles of 61-64^{∘}, although with a wider range of H–C⋯C angles (ca. 60^{∘} to 75^{∘}). Results for all members of the series will be presented and compared to theoretical data.

aXIAM v2.5e. H. Hartwig and H. Dreizler, Z. Naturforsch \underline{51a}, (1996), 923.